# Adaptive Modality Balanced Online Knowledge Distillation for Brain-Eye-Computer based Dim Object Detection

### v1

@@ I Introduction
Figure 1: Data extraction and processing procedure based on adaptive modality balanced online knowledge distillation (AMBOKD) method.
@@
This paper is structured as follows: Section. II discusses the related work about this study. Section. III presents the design of the built system and the ESSVP dataset in detail. Section. IV illustrates the proposed target recognition method. Section. V presents the experimental results and analyzes the performance. Section. VI draws conclusions and possible directions for future work.
@@ II-B Multimodal Learning
Modeling and analyzing data from various sensory modalities, including images, speech, text, and EEG signals, are key aspects of multimodal learning [32, 33, 34]. Significant attention has been garnered by multimodal learning in the fields of artificial intelligence and machine learning in recent years, resulting in remarkable progress [9, 35]. Currently, different multimodal applications, such as human-computer interaction [33, 36], natural language translation [37, 38], and computer vision [39, 40, 34], are being applied in our lives.
Due to the development of deep learning, researchers propose novel loss functions or training approaches to optimize multimodal models. For instance, Nagrani et al. [41] introduced a transformer-based architecture that employs fusion bottlenecks for modality fusion at multiple layers. This method enables each modality to capture crucial information while efficiently sharing necessary information, leading to enhanced fusion performance and reduced computational cost. Du et al. [32] modeled the relationships between brain, visual, and linguistic features using multimodal deep generative models. They maximized both intra-modality and inter-modality mutual information regularization terms. Their approach addresses limitations such as the under-exploitation of multimodal knowledge and the scarcity of training data. Yang et al. [40] proposed a multimodal fusion approach for remote sensing image-audio retrieval tasks. They converted audio inputs to text and fused them with the text information to obtain a fusion representation. They optimized the common retrieval space using triplet loss, semantic loss, and consistency loss. Experimental findings on multiple datasets indicate the effectiveness of their approach.
Most of the current methods in this field primarily focus on feature fusion and loss function design when developing multimodal fusion algorithms, thereby overlooking the potential benefits of knowledge transfer and mutual learning between different modalities. In contrast, we propose a dynamic hybrid fusion approach that integrates multimodal features and dynamically trains all branches in the model using an OKD method. This mutual learning approach for multimodal fusion aims to fully leverage the benefits of the visual and cognitive domains and extract crucial information effectively.
@@ II-C Knowledge Distillation
In the field of multimodal learning, attention has been drawn to offline knowledge distillation due to its knowledge transfer capabilities. For instance, Zhang et al. [48] proposed a visual-to-EEG cross-modal KD method that enhances continuous EEG prediction using dark knowledge from visual modality. Multi-teacher KD methods, including AvgMKD [49], CA-MKD [50], and EMKD [51], have been proposed to combine knowledge from various modalities and train the student model.
Current online knowledge distillation approaches have achieved significant progress with multimodal homogeneous data. For instance, Zhang et al. [47] introduced DML, demonstrating that student models can learn from each other through their predictions in deep mutual learning. Guo et al.[52] presented an OKD approach using collaborative learning with a weighted ensemble logit distribution. Li et al. [53] proposed an innovative embedded OKD approach that surpasses existing OKD approaches in image classification tasks. This approach leverages ensemble information, overall feature representations from peer networks, and logits to fully exploit the potential of networks.
However, there remains a research gap in online knowledge distillation methods for multimodal heterogeneous data, due to the challenges in extracting and fusing heterogeneous data features. Thus, we propose the AMBOKD method, which fully extracts features from both 2D EEG data and 3D image data, and treats the features output by the fusion model as a single modality for mutual learning. This method adaptively adjust the influences weights and the optimization level of each modality, leading to fully optimization of the parameters and surperior performance of the fusion model, thereby opening up new application scenarios for multimodal OKD.
@@ III-B Suspicious Region Detection
As illustrated in Fig. 2, the designed system uses a pretrained ResNet50 feature encoder to obtain feature maps of image data, and employs RPN to generate region proposals, which are denoted with black boxes. The anchors used have scales of 16, 32, 48, 64, and 80, with aspect ratios of 0.5, 0.75, 1.0, 1.5, and 2.0. We employ UAVs to capture images of toy models in open outdoor settings, forming the training set. Meanwhile, the test set comprises images taken in complex real-world environments. The effectiveness of the employed feature extraction method and the RPN network is validated by the results, as depicted in Fig. 3. It is evident that, despite variations in scene environments and targets, the RPN network is capable of precisely delineating suspicious objects, thereby effectively supporting the continuation of subsequent experiments.
@@ III-D Data Collection and Preprocessing
This dataset consists of 13405 samples, including 3880 positive samples with the target image and 9525 negative samples without the target image. Each sample includes EEG data from 59 electrode channels, as 5 out of the 64 electrodes are redundant and therefore excluded. The EEG data are filtered using a 2$\sim$30 Hz bandpass filter and downsampled from the original 1000 Hz to a rate of 250 Hz. The data is then segmented into 1.2 s (-500$\sim$700 ms) samples based on the trigger. Baseline correction is applied by employing the data from the first 200 ms interval as a reference to subtract the mean activity level. The image data used has a minimum pixel size of $150\times 150$, corresponding to the center of the focused candidate box. All images are resized to $224\times 224$ pixels to standardize the input for image network. The images in the validation set are preprocessed by introducing noise to simulate real-world conditions and evaluate the algorithm’s robustness, as shown in Fig. 6. Specifically, one-third of the images are augmented with 0.2 Gaussian noise and another third of the images are augmented with 0.2 salt and pepper noise.
@@ IV Methodology
The framework of adaptive modality balanced online knowledge distillation (AMBOKD) method is illustrated in Fig. 7. Initially, AMBOKD receives paired data $X$ and extract EEG and visual features through the visual encoder and EEG encoder. These features are subsequently passed through the multi-head self-attention module to generate the fusion features, and the features from each domain are sent to their respective classifiers. The features extracted from the visual, EEG, and fusion domains are classified and employed to compute the losses of the three modalities using the cross-entropy (CE) and KD loss functions. Subsequently, the three modalities take turns as students and teachers to compute the total loss, the weight of the teacher influence, and the dynamic student gradient for parameter updating. The processes of the feature extraction, feature fusion, the online knowledge distillation, and the adaptive modality balancing are detailed as follows.
IV-A Extraction of The Modality Representation
@@ IV-A1 Visual Encoder
The effective representation of visual images from the visual encoder can be easily extracted with the help of EefficientNet. The representation of the visual domain is computed as follows:
@@ IV-A2 EEG Encoder
In summary, MCGRAM can capture and encode relevant information from EEG signals by cascading frequency, spatial, and temporal blocks. This resulted in a comprehensive and informative representation for further analysis and classification tasks. The representation of the EEG modality is computed as:
@@ IV-D1 Dynamic Weights for KD Losses
Algorithm 1 Dynamic weights modulation process
0:  Training Dataset $X$, iteration number $T$, batchsize $B_{m}$, knowledge distillation temperature $\tau$
  for $t=0,1,\ldots,T$ do
[…]
     Feed-forward the batched data $B_{t}$ to the model
     Compute the CE losses $\mathcal{L}_{CE_{(t)}}^{S}$, $\mathcal{L}_{CE_{(t)}}^{T_{a}}$ and $\mathcal{L}_{CE_{(t)}}^{T_{b}}$
     Compute the KD losses $\mathcal{L}_{KD_{(t)}}^{T_{a}}$ and $\mathcal{L}_{KD_{(t)}}^{T_{b}}$
     Compute $\alpha=\text{Sat}\bigg{(}\frac{\mathcal{L}_{CE_{(t)}}^{S}}{\mathcal{L}_{CE_{(t)}}^{T_{a}}},{\alpha}_{\min},{\alpha}_{\max}\bigg{)}$
     Compute $\beta=\text{Sat}\bigg{(}\frac{\mathcal{L}_{CE_{(t)}}^{S}}{\mathcal{L}_{CE_{(t)}}^{T_{b}}},{\beta}_{\min},{\beta}_{\max}\bigg{)}$
     Compute the total loss ${L}_{{total}_{(t)}}^{S}=\mathcal{L}_{CE_{(t)}}^{S}+\alpha{\tau^{2}}\mathcal{L}_{KD_{(t)}}^{T_{a}}+\beta{\tau^{2}}\mathcal{L}_{KD_{(t)}}^{T_{b}}$
  end for
@@ IV-D2 Dynamic Ratios for Backward Gradients
The updating process is presented in Algorithm 2. Specifically, the ratios $R^{S}_{(t)}$, $R^{T_{a}}_{(t)}$ and $R^{T_{b}}_{(t)}$ are computed for indicating the current training optimization level of each modality. To ensure consistent training optimization progress and maximize the completeness and effectiveness of the optimization in multimodal mutual learning, the dynamic learning progress ratio in $t$-th iteration step is computed as:
$R_{(t)}^{DG}=\begin{cases}1&\text{if }E_{n}=1\\ \text{Sat}\left(\big{(}\frac{R_{(t)}^{T_{a}}+R_{(t)}^{T_{b}}}{2\times R_{(t)}^{S}}\big{)}^{\gamma},R_{\min},R_{\max}\right)&\text{if }E_{n}>1\end{cases}$ (13)
@@
${\theta}^{S}_{(t+1)}\leftarrow{\theta}^{S}_{(t)}-R_{(t)}^{DG}\frac{\eta}{\sqrt{D^{\rm{var}}_{(t)}}+\epsilon}\cdot D^{\rm{mean}}_{(t)}$ (14)
@@
Algorithm 2 Dynamic gradients modulation process
  for $t=0,1,\ldots,T$ do
[…]
     Feed-forward the batched data $B_{t}$ to the model
     Compute the losses $\mathcal{L}_{CE_{(t)}}^{S}$, $\mathcal{L}_{CE_{(t)}}^{T_{a}}$, $\mathcal{L}_{CE_{(t)}}^{T_{b}}$ and $\ell_{total}\left(x;\theta_{(t)}^{S}\right)$
     if $E_{n}==1$ then
        Set $R_{(t)}^{DG}\leftarrow 1$
     else
        Update $R^{S}_{(t)}\leftarrow\frac{\mathcal{L}_{\rm{base}}^{S}-\mathcal{L}_{CE_{(t)}}^{S}}{\mathcal{L}_{\rm{base}}^{S}}$, $R_{(t)}^{T_{a}}\leftarrow\frac{\mathcal{L}_{\rm{base}}^{T_{a}}-\mathcal{L}_{CE_{(t)}}^{T_{a}}}{\mathcal{L}_{\rm{base}}^{T_{a}}}$, $R_{(t)}^{T_{b}}\leftarrow\frac{\mathcal{L}_{\rm{base}}^{T_{b}}-\mathcal{L}_{CE_{(t)}}^{T_{b}}}{\mathcal{L}_{\rm{base}}^{T_{b}}}$
        Compute $R_{(t)}^{DG}=\text{Sat}\big{(}\big{(}\frac{R_{(t)}^{T_{a}}+R_{(t)}^{T_{b}}}{2\times R_{(t)}^{S}}\big{)}^{\gamma},R_{\min},R_{\max}\big{)}$
     end if
     Compute $\tilde{g}\left(\theta_{(t)}^{S}\right)=\frac{1}{B_{m}}\sum_{x\in B_{(t)}}\nabla_{\theta^{S}}\ell_{total}\left(x;\theta_{(t)}^{S}\right)$
     Compute $D^{\rm{mean}}_{(t)}=\beta_{1}\cdot D^{\rm{mean}}_{(t-1)}+\left(1-\beta_{1}\right)\cdot\tilde{g}\big{(}\theta_{(t)}^{S}\big{)}$
     Compute $D^{\rm{var}}_{(t)}=\beta_{2}\cdot D^{\rm{var}}_{(t-1)}+\left(1-\beta_{2}\right)\cdot\tilde{g}^{2}\big{(}\theta_{(t)}^{S}\big{)}$
     Update ${\theta}^{S}_{(t+1)}\leftarrow{\theta}^{S}_{(t)}-R_{(t)}^{DG}\frac{\eta}{\sqrt{D^{\rm{var}}_{(t)}}+\epsilon}\cdot D^{\rm{mean}}_{(t)}$
  end for
  if $E_{n}==1$ then
     Calculate average CE losses $\mathcal{L}_{\rm{base}}^{S},\mathcal{L}_{\rm{base}}^{T_{a}},\mathcal{L}_{\rm{base}}^{T_{b}}$
  end if
@@ V-C2 Visualization Analysis
The latent representations of the models in the ablation analysis are visualized using the t-distributed stochastic neighbor embedding (t-SNE) approach to further verify the feature extraction and fusion capability of our proposed approach. As depicted in Fig. 10, the feature distributions of all subjects are displayed with various colors of dots, with red dots representing the features in the target domain, and dots in other colors representing the features in the nontarget class.
Specifically, unimodality models such as EfficientNet and MCGRAM exhibit distinct visual and EEG modality characteristics in their feature representations, as illustrated in Fig. 10. The visual-based EfficientNet model exhibits a more discrete feature representation, with a large intra-class distance, indicating its superior feature extraction performance but poor generalization performance of the classifier. Conversely, the feature representation of the MCGRAM model based on EEG is more clustered, with a small intra-class distance, indicating weaker feature extraction ability but stronger classifier generalization. As a multimodal fusion model, AMM effectively combines the benefits of vision and EEG modalities, thereby further reducing the inter-class distance while maintaining the feature representation capability. By integrating single-teacher guidance mechanisms into the multimodal fusion model, V-KD and E-KD models exhibit enhanced classification performance and generalization, as reflected in larger inter-class and smaller intra-class differences in feature representation. In addition, by introducing a multi-teacher mechanism, OKD approach, and multimodal AMB approach, the discrimination of feature representation extracted by MKD, MMOKD, and AMBOKD models become more pronounced and the feature representation distances within the same class become closer. This indicates that these mechanisms enhance the capability of multimodal fusion and effectively improve the classifier’s accuracy and generalization.

### v2

@@ I Introduction
Figure 1: Data extraction and processing procedure based on adaptive modality balanced online knowledge distillation (AMBOKD) method. See https://youtu.be/Yj01V8TBlUQ for an overview of the experiment and the AMBOKD method.
@@
The reset of this paper is structured as follows: Section II discusses the related work about this study. Section III presents the design of the built system and the ESSVP dataset in detail. Section IV illustrates the proposed target recognition method. Section V presents the experimental results and analyzes the performance. Section VI draws conclusions and possible directions for future work.
@@ II-B Multimodal Learning
Modeling and analyzing data from various sensory modalities, including images, speech, text, and EEG signals, are key aspects of multimodal learning [32, 33, 34]. Significant attention has been garnered by multimodal learning in the fields of artificial intelligence and machine learning in recent years, resulting in remarkable progress [9, 35]. Currently, different multimodal applications, such as human-computer interaction [33, 36], natural language translation [37, 38], and computer vision [39, 34], are being applied in our lives.
Due to the development of deep learning, researchers propose novel loss functions or training approaches to optimize multimodal models. For instance, Nagrani et al. [40] introduced a transformer-based architecture that employs fusion bottlenecks for modality fusion at multiple layers. This method enables each modality to capture crucial information while efficiently sharing necessary information, leading to enhanced fusion performance and reduced computational cost. Du et al. [32] modeled the relationships between brain, visual, and linguistic features using multimodal deep generative models. They maximized both intra-modality and inter-modality mutual information regularization terms. Their approach addresses limitations such as the under-exploitation of multimodal knowledge and the scarcity of training data. Yang et al. [41] proposed a multimodal fusion approach for remote sensing image-audio retrieval tasks. They converted audio inputs to text and fused them with the text information to obtain a fusion representation. They optimized the common retrieval space using triplet loss, semantic loss, and consistency loss. Experimental findings on multiple datasets indicate the effectiveness of their approach. However, these methods primarily focus on feature fusion and loss function design when developing multimodal fusion algorithms, thereby overlooking the potential benefits of knowledge transfer and mutual learning between different modalities.
In this paper, we propose a dynamic hybrid fusion approach that integrates multimodal features and dynamically trains all branches in the model using an OKD method. This mutual learning approach for multimodal fusion aims to fully leverage the benefits of the visual and cognitive domains and extract crucial information effectively.
@@ II-C Knowledge Distillation
In the field of multimodal learning, attention has been drawn to offline knowledge distillation due to its knowledge transfer capabilities. Zhang et al. [48] proposed a visual-to-EEG cross-modal KD method that enhances continuous EEG prediction using dark knowledge from visual modality. Multi-teacher KD methods, including AvgMKD [49], CA-MKD [50], and EMKD [51], have been proposed to combine knowledge from various modalities and train the student model.
Current online knowledge distillation approaches have achieved significant progress with multimodal homogeneous data. For instance, Zhang et al. [47] introduced DML, demonstrating that student models can learn from each other through their predictions in deep mutual learning. Guo et al.[52] presented an OKD approach using collaborative learning with a weighted ensemble logit distribution. Li et al. [53] proposed an innovative embedded OKD approach that surpasses existing OKD approaches in image classification tasks. This approach leverages ensemble information, overall feature representations from peer networks, and logits to fully exploit the potential of networks. However, there remains a research gap in online knowledge distillation methods for multimodal heterogeneous data, due to the challenges in extracting and fusing heterogeneous data features.
To this end, we propose the AMBOKD method, which fully extracts features from both 2D EEG data and 3D image data, and treats the features output by the fusion model as a single modality for mutual learning. This method adaptively adjust the influences weights and the optimization level of each modality, leading to fully optimization of the parameters and surperior performance of the fusion model, thereby opening up new application scenarios for multimodal OKD.
@@ III System Design and Data Collection
In this section, we start by introducing the structure and data acquisition process of the brain-eye-computer based object detection system. Then, we verify the effectiveness of the employed RPN network. The designed ESSVP paradigm and its experimental environment are illustrated in detail. Lastly, we present the preprocessing process and the details of the ESSVP dataset.
@@ III-B Suspicious Region Detection
As illustrated in Fig. 2, the designed system uses a pretrained ResNet50 feature encoder to obtain feature maps of image data, and employs RPN to generate region proposals, which are denoted with black boxes. The anchors used have scales of 16, 32, 48, 64, and 80, with aspect ratios of 0.5, 0.75, 1.0, 1.5, and 2.0. We employ UAVs to capture images of toy models in open outdoor settings, forming the training set. Meanwhile, the test set comprises images taken in complex real-world environments. Fig. 3 shows the difference between the training set and the validation set. It can be seen that although the background of the validation set is complex and the target size is small, the RPN network trained by the training set can still effectively detect suspicious targets. These results confirm the effectiveness of the employed RPN network.
@@ III-D Data Collection and Preprocessing
This dataset consists of 13405 samples, including 3880 positive samples with the target image and 9525 negative samples without the target image. Each sample includes EEG data from 59 electrode channels, as 5 out of the 64 electrodes are redundant and therefore excluded. The EEG data are filtered using a 2$\sim$30 Hz bandpass filter and downsampled from the original 1000 Hz to a rate of 250 Hz. The data is then segmented into 1.2 s (-500$\sim$700 ms) samples based on the trigger. Baseline correction is applied by employing the data from the first 200 ms interval as a reference to subtract the mean activity level. The image data used has a minimum pixel size of $150\times 150$, corresponding to the center of the focused candidate box. All images are resized to $224\times 224$ pixels to standardize the input for image network. The images in the validation set are preprocessed by introducing noise to simulate real-world conditions and evaluate the algorithm’s robustness, as shown in Fig. 6. Specifically, one-third of the images are enhenced with 0.2 Gaussian noise and another third of the images are enhenced with 0.2 salt and pepper noise.
@@ IV Methodology
The framework of the proposed adaptive modality balanced online knowledge distillation (AMBOKD) method is illustrated in Fig. 7. Initially, AMBOKD receives paired data $X$ and extract EEG and visual features through the visual encoder and EEG encoder. These features are subsequently passed through the multi-head self-attention module to generate the fusion features, and the features from each domain are sent to their respective classifiers. The features extracted from the visual, EEG, and fusion domains are classified and employed to compute the losses of the three modalities using the cross-entropy (CE) and KD loss functions. Subsequently, the three modalities take turns as students and teachers to compute the total loss, the weight of the teacher influence, and the dynamic student gradient for parameter updating. The processes of the feature extraction, feature fusion, the online knowledge distillation, and the adaptive modality balancing are detailed as follows.
IV-A Extraction of Modality Representation
@@ IV-A1 Visual Encoder
The effective representation of visual images from the visual encoder can be easily extracted with the help of EefficientNet. The representation of the visual domain is computed as
@@ IV-A2 EEG Encoder
In summary, MCGRAM can capture and encode relevant information from EEG signals by cascading frequency, spatial, and temporal blocks. This resulted in a comprehensive and informative representation for further analysis and classification tasks. The representation of the EEG modality is computed as
@@ IV-D2 Dynamic Ratios for Backward Gradients
Algorithm 1 The parameters updating process
1:  for $t=0,1,\ldots,T$ do
[…]
3:     Feed-forward the batched data $B_{t}$ to the model /** Dynamic weights modulation **/
4:     Compute the CE losses $\mathcal{L}_{CE_{(t)}}^{S}$, $\mathcal{L}_{CE_{(t)}}^{T_{a}}$ and $\mathcal{L}_{CE_{(t)}}^{T_{b}}$
5:     Compute the KD losses $\mathcal{L}_{KD_{(t)}}^{T_{a}}$ and $\mathcal{L}_{KD_{(t)}}^{T_{b}}$
6:     Compute $\alpha=\text{Sat}\Big{(}\frac{\mathcal{L}_{CE_{(t)}}^{S}}{\mathcal{L}_{CE_{(t)}}^{T_{a}}},{\alpha}_{\min},{\alpha}_{\max}\Big{)}$
7:     Compute $\beta=\text{Sat}\Big{(}\frac{\mathcal{L}_{CE_{(t)}}^{S}}{\mathcal{L}_{CE_{(t)}}^{T_{b}}},{\beta}_{\min},{\beta}_{\max}\Big{)}$
8:     Compute the total loss $\mathcal{L}_{{total}_{(t)}}^{S}=\mathcal{L}_{CE_{(t)}}^{S}+\alpha{\tau^{2}}\mathcal{L}_{KD_{(t)}}^{T_{a}}+\beta{\tau^{2}}\mathcal{L}_{KD_{(t)}}^{T_{b}}$ /** Dynamic gradients modulation **/
9:     if $E_{n}==1$ then
10:        Calculate average CE losses $\mathcal{L}_{\rm{base}}^{S},\mathcal{L}_{\rm{base}}^{T_{a}},\mathcal{L}_{\rm{base}}^{T_{b}}$
11:        Set $R_{(t)}^{DG}\leftarrow 1$
12:     else
13:        Update $R^{S}_{(t)}\leftarrow\frac{\mathcal{L}_{\rm{base}}^{S}-\mathcal{L}_{CE_{(t)}}^{S}}{\mathcal{L}_{\rm{base}}^{S}}$, $R_{(t)}^{T_{a}}\leftarrow\frac{\mathcal{L}_{\rm{base}}^{T_{a}}-\mathcal{L}_{CE_{(t)}}^{T_{a}}}{\mathcal{L}_{\rm{base}}^{T_{a}}}$, $R_{(t)}^{T_{b}}\leftarrow\frac{\mathcal{L}_{\rm{base}}^{T_{b}}-\mathcal{L}_{CE_{(t)}}^{T_{b}}}{\mathcal{L}_{\rm{base}}^{T_{b}}}$
14:        Compute $R_{(t)}^{DG}=\text{Sat}\Big{(}\big{(}\frac{R_{(t)}^{T_{a}}+R_{(t)}^{T_{b}}}{2\times R_{(t)}^{S}}\big{)}^{\gamma},R_{\min},R_{\max}\Big{)}$
15:     end if
16:     Compute $\tilde{g}\left(\theta_{(t)}^{S}\right)=\frac{1}{B_{m}}\sum_{x\in B_{(t)}}\nabla_{\theta^{S}}\mathcal{L}_{{total}_{(t)}}^{S}$
17:     Compute $D^{\rm{mean}}_{(t)}=\beta_{1}\cdot D^{\rm{mean}}_{(t-1)}+\left(1-\beta_{1}\right)\cdot\tilde{g}\big{(}\theta_{(t)}^{S}\big{)}$
18:     Compute $D^{\rm{var}}_{(t)}=\beta_{2}\cdot D^{\rm{var}}_{(t-1)}+\left(1-\beta_{2}\right)\cdot\tilde{g}^{2}\big{(}\theta_{(t)}^{S}\big{)}$
19:     Update ${\theta}^{S}_{(t+1)}\leftarrow{\theta}^{S}_{(t)}-R_{(t)}^{DG}\cdot\frac{\eta}{\sqrt{D^{\rm{var}}_{(t)}}+\epsilon}\cdot D^{\rm{mean}}_{(t)}$
20:  end for
20:  ${\theta}^{S}$
@@
The updating process is presented in Algorithm 1. Specifically, the ratios $R^{S}_{(t)}$, $R^{T_{a}}_{(t)}$ and $R^{T_{b}}_{(t)}$ are computed for indicating the current training optimization level of each modality. To ensure consistent training optimization progress and maximize the completeness and effectiveness of the optimization in multimodal mutual learning, the dynamic learning progress ratio in $t$-th iteration step is computed as
$R_{(t)}^{DG}=\begin{cases}1&\text{if }E_{n}=1\\ \text{Sat}\Big{(}\big{(}\frac{R_{(t)}^{T_{a}}+R_{(t)}^{T_{b}}}{2\times R_{(t)}^{S}}\big{)}^{\gamma},R_{\min},R_{\max}\Big{)}&\text{if }E_{n}>1\end{cases}$ (13)
@@
${\theta}^{S}_{(t+1)}\leftarrow{\theta}^{S}_{(t)}-R_{(t)}^{DG}\frac{\eta}{\sqrt{D^{\rm{var}}_{(t)}}+\epsilon}\cdot D^{\rm{mean}}_{(t)},$ (14)
@@ V-C2 Visualization Analysis
The latent representations of the models in the ablation analysis are visualized using the t-distributed stochastic neighbor embedding (t-SNE) approach to further verify the feature extraction and fusion capability of our proposed approach [68]. As depicted in Fig. 10, the feature distributions of all subjects are displayed by different colored dots, with red dots representing the features in the target domain, and dots in other colors representing the features in the non-target class.
As can be seen from Fig. 10, unimodality models such as EfficientNet and MCGRAM exhibit distinct visual and EEG modality characteristics in their feature representations. The visual-based EfficientNet model exhibits a more discrete feature representation, with a large intra-class distance, indicating its superior feature extraction performance but poor generalization performance of the classifier. Conversely, the feature representation of the MCGRAM model based on EEG is more clustered, with a small intra-class distance, indicating weaker feature extraction ability but stronger classifier generalization. As a multimodal fusion model, AMM effectively combines the benefits of vision and EEG modalities, thereby further reducing the inter-class distance while maintaining the feature representation capability. By integrating single-teacher guidance mechanisms into the multimodal fusion model, V-KD and E-KD models exhibit enhanced classification performance and generalization, as reflected in larger inter-class and smaller intra-class differences in feature representation. In addition, by introducing a multi-teacher mechanism, OKD approach, and multimodal AMB approach, the discrimination of feature representation extracted by MKD, MMOKD, and AMBOKD models become more pronounced and the feature representation distances within the same class become closer. This indicates that these mechanisms enhance the capability of multimodal fusion and effectively improve the classifier’s accuracy and generalization.